OpenStudy (anonymous):

Solve: (-1to 0)∫[(3t)/(4+t^2)^4][dt] NOTE:(By-1to 0 I mean that the -1 is below the integral and the 0 is up or above the integral)

zepdrix (zepdrix):

\[\huge \int\limits_{-1}^0 \frac{3t}{(4+t^2)^4}dt\]We'll want to apply a u substitution. When looking for what to substitute, it's fairly simple with polynomials. See how we have a t^2 and a t^1 in this problem? If we let our t^2 term be U, then we take the derivative of U, we'll get something involving t^1 for our DU.

OpenStudy (anonymous):

wouldn't DU=2t...??

zepdrix (zepdrix):

Yes it would! And our top is 3t, hmmm that's no good. But it's not a problem, we can work with it. We just need them to be of the same DEGREE (power).

zepdrix (zepdrix):

\[\huge \text{Let} \quad u=4+t^2 \; \text{, then} \quad du=2t \cdot dt\]

zepdrix (zepdrix):

Now the idea is, we want to try and solve for t dt. Let me rearrange the top of the equation so you can see what I'm talking about.\[\huge 3\int\limits\limits_{-1}^0 \frac{(t\cdot dt)}{(4+t^2)^4}\]

zepdrix (zepdrix):

So in our du, let's divide both sides by 2, so we have it written in terms of t*dt.\[\huge \frac{1}{2}du=t \cdot dt\]

zepdrix (zepdrix):

Now we can successfully replace all of our T and DT with U DU.

zepdrix (zepdrix):

\[\huge 3\int\limits\limits\limits_{-1}^0 \frac{(t\cdot dt)}{(4+t^2)^4} \qquad \rightarrow \qquad \huge 3\int\limits\limits\limits_{-1}^0 \frac{(\frac{1}{2}du)}{(u)^4}\]

zepdrix (zepdrix):

Any part of that have you stumped? It's a bit tricky getting used to substitutions.

OpenStudy (anonymous):

SO far its good. You haven't lost me yet, I was doing it out as you were explaining...

zepdrix (zepdrix):

\[\huge \frac{3}{2}\int\limits\limits\limits\limits_{-1}^0 \frac{1}{u^4}du \qquad \rightarrow \qquad \huge \frac{3}{2}\int\limits\limits\limits\limits_{-1}^0 u^{-4}du\] And from here we're able to apply the Power Rule for Integration.

OpenStudy (anonymous):

ok so it'd be[ 4^-3/-3] +c ...do i have to multiply by the (3/2)??

zepdrix (zepdrix):

why does your U look like a 4? lol Wait this feels very familiar, -_- someone else was having that problem before also XD Is it a strange keyboard or something? :3

OpenStudy (anonymous):

it does ?? 0_0 T.T

zepdrix (zepdrix):

\[\huge \frac{3}{2}\cdot \left(-\frac{1}{3}u^{-3}\right)\]Yes you have the right idea.

OpenStudy (anonymous):

ohh lolol T.T

OpenStudy (anonymous):

i messed up T.T you are too kind

zepdrix (zepdrix):

Now we have to be careful.... The limits of integration were for our variable T! Not U. So from here, we have 2 options. We can either... Find new limits of integration, plugging our limits into the U sub. Or Back substitute, plug our (4+t^2) back in for U and then apply the limits of integration that we started with. They'll both take about the same amount of work to apply, so whichever method you prefer.

OpenStudy (anonymous):

find new limits of integration, the first one pls.

zepdrix (zepdrix):

\[\huge u=4+t^2 \]Plugging in our previous limits of integration gives us,\[\large t=-1 \qquad \rightarrow \qquad u=4+(-1)^2 \qquad \rightarrow \qquad u=5\]\[\large t=0 \qquad \rightarrow \qquad u=4+(0)^2 \qquad \rightarrow \qquad u=4\]

zepdrix (zepdrix):

So now our integral goes from 5 to 4.

zepdrix (zepdrix):

Cancelling the 3's out from our U solution gives us...\[\huge \frac{3}{2}\cdot \left(-\frac{1}{3}u^{-3}\right)|_4^5 \qquad \rightarrow \qquad -\frac{1}{2}\left(4^{-3}-5^{-3}\right)\]

zepdrix (zepdrix):

Do you have an answer key or anything, so we can check our work? D':

OpenStudy (anonymous):

ohh hahaha i was feeling dumb bc i still couldn't figure out how I came up with the 4, what happened was that when I wrote it down on my paper it my 'u" looked like a 4 T.T

OpenStudy (anonymous):

Unfortunately, no. But I will double check with my prof tomorrow.

OpenStudy (anonymous):

Thank you so much!! ^_^

zepdrix (zepdrix):

Hopefully I didn't make a mistake anywhere in there c: Yay team \:D/